OpenStudy (anonymous):

the distribution of body weights has a mean of 188 pounds and a standard deviation of 7. If 267 men belong to the gym, how many of them do you expect to be over 200 pounds?

OpenStudy (anonymous):

Since 188 lbs is mean of 267 men in the gym, z = (200 - 188 ) / 7 = 1.714 P ( x > 200 ) = P ( z > 1.714 )

OpenStudy (anonymous):

AH! THANK YOU!

OpenStudy (anonymous):

wait here are the answers they give to chose from About 12 or 13 men B About 10 or 11 men C About 11 or 12 men D About 255 or 256 men

OpenStudy (anonymous):

You need the table to look up Z score!

OpenStudy (anonymous):

Hold on!

OpenStudy (anonymous):

okay!

OpenStudy (anonymous):

still with me?

OpenStudy (anonymous):

I can't find the table, you probably has it in your book appendix!

OpenStudy (anonymous):

I'm not sure how to use the table..

OpenStudy (anonymous):

Match the row has number "1" with the column .71

OpenStudy (anonymous):

.9591

OpenStudy (anonymous):

i think..

OpenStudy (anonymous):

wait thats wrong

OpenStudy (anonymous):

.9649

OpenStudy (anonymous):

Are you sure?

OpenStudy (anonymous):

pretty sure..

OpenStudy (anonymous):

i went to the 1.7 row, then the .01 column...that's what my z table had

OpenStudy (anonymous):

So now you know how to use it, right?

OpenStudy (anonymous):

well, idk what i'm supposed to do with tha number, to get one of the answers i gave you above..

OpenStudy (anonymous):

Multiply .9649 * 267 = 257 men

OpenStudy (anonymous):

ooo! i was trying to do it to 188 men. lol

OpenStudy (anonymous):

About that number!

OpenStudy (anonymous):

thanks for the help!

OpenStudy (anonymous):

I just don't have the table to match up, but I trust you read it right!

OpenStudy (anonymous):

i'm about to post another question